1.Name of the Employee Abebe Gemeda Basic Salary 12,000 Transport

Allowance 4,000 HardshipAllowance 3000 over time Hours 16 Duration Holiday.

2. Name of the Employee Ayisha Sultan Basic Salary 6000 Transport Allowance 3000 Hardship Allowance 2000 Over time Hours 12 Duration normal.

3.Name of the Employee Chala Tola Basic Salary 8,000 Transport Allowance 3500 Hardship Allowance 2500 Over time Hours 10 Duration weekend.

4.Name of the Employee Ahmed Seid Basic Salary 5000 ransport Allowance 2500 Hardship Allowance 1500 Over time Hours 8 Duration late .

Additional Information

1. All employees are expected to work for 40 hours per week and a month has 4 weeks. And all employees perform as it is expected

2. All employees except Chala Tola are permanent

3. All employees except Ayisha Sultan Contribute 4% to wards credit association

4. The employer (ABC Company) contributes 20% of each permanent employee’s towards provident fund

Based on the data given above calculate

A. Overtime of each employee



1
Expert's answer
2021-07-01T05:04:13-0400

Abebe Gemeda:

"\frac{12000}{160}*16=1200"

Ayisha Sultan

"\frac{6000}{160}*12=450"

Chala Tola 

"\frac{8000}{160}*10=500"

Ahmed Seid

"\frac{5000}{160}*8=250"